# What You See Is Not Always What You Get: An Empirical Study of Code Comprehension by Large Language Models

### v1

@@ Abstract.
Recent studies have demonstrated outstanding capabilities of large language models (LLMs) in software engineering domain, covering numerous tasks such as code generation and comprehension. While the benefit of LLMs for coding task is well noted, it is perceived that LLMs are vulnerable to adversarial attacks. In this paper, we study the specific LLM vulnerability to imperceptible character attacks, a type of prompt-injection attack that uses special characters to befuddle an LLM whilst keeping the attack hidden to human eyes. We devise four categories of attacks and investigate their effects on the performance outcomes of tasks relating to code analysis and code comprehension. Two generations of ChatGPT are included to evaluate the impact of advancements made to contemporary models. Our experimental design consisted of comparing perturbed and unperturbed code snippets and evaluating two performance outcomes, which are model confidence using log probabilities of response, and correctness of response. We conclude that earlier version of ChatGPT exhibits a strong negative linear correlation between the amount of perturbation and the performance outcomes, while the recent ChatGPT presents a strong negative correlation between the presence of perturbation and performance outcomes, but no valid correlational relationship between perturbation budget and performance outcomes. We anticipate this work contributes to an in-depth understanding of leveraging LLMs for coding tasks. It is suggested future research should delve into how to create LLMs that can return a correct response even if the prompt exhibits perturbations.
Imperceptible character attack, Large language models, Code comprehension
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligenceccs: Software and its engineeringccs: Software and its engineering General programming languages
@@ 1. Introduction
The effectiveness of Large Language Models (LLMs) at natural language processing has led to their popularity and application across many different fields, one of which is as an assistive tool for software developers. The ability to interact with them using natural language makes it very easy for developers to integrate such tools into their existing workflows. LLMs can assist with tasks such as code generation (Du et al., 2024), program repair (Zhang et al., 2023), and vulnerability detection (Zhou et al., 2024).
However, appearing to respond like a human does not mean that an LLM can reason the way a human can, and integrating these systems into workflows means taking on new vulnerabilities brought on by these LLMs. Considering their widespread adoption and continued growth in their sector, it is important to ensure the robustness of LLMs with regards to natural language processing tasks through studying their responses to attacks and understanding their mechanisms.
@@
There have been several recent studies on text-based attacks against LLMs, targeting performance outcomes of natural language processing (NLP) tasks (Boucher et al., 2022)(Formento et al., 2021)(Li et al., 2020). While these studies have been successful in proving the feasibility of text-based attacks, the feasibility of these attacks being applied in a real-world setting is still in question. In (Dyrmishi et al., 2023), it is found that an effective adversarial attack against a text-based LLM should produce text that is both valid and natural. Attacks that do not exhibit these characteristics may cause failed NLP model decisions, but their real-world effectiveness is limited if semantic consistency is not maintained, or if the adversarial example is clearly computer-generated. Through a survey of 378 human participants, Dyrmishi et al. concludes that existing text attacks are impractical in real-world scenarios where humans are involved (Dyrmishi et al., 2023).
To address these issues, this paper presents an in-depth empirical study to determine the efficacy of a truly imperceptible text-based attack against LLMs. We define ‘true imperceptibility’ as a malicious prompt that is completely visually indistinguishable from a benign prompt. An ‘imperceptible’ attack such as the one outlined in (Liu et al., 2021) is not considered as truly imperceptible, as the changed prompt is visually distinguishable from the benign prompt; though steps have been taken to make them visually and semantically similar, the prompts are clearly different when placed side-by-side.
@@
To maintain true imperceptibility, the attack method studied creates perturbed prompts by adding special Unicode specification characters that affect visual rendering to clean prompts. Taking inspiration from Boucher et al. (Boucher et al., 2022), four categories of imperceptible attack are delineated: reordering, invisible characters, deletions, and homoglyphs. Following this, we devise four types of imperceptible adversarial attacks for code analysis and comprehension task. Attacks are performed against three different ChatGPT models: two third-generation models and one fourth-generation model. The efficacy of attacks are determined based on their effect on the models’ performance outcomes of confidence and correctness in a code analysis task. Experiment results demonstrates a strong negative correlation between perturbation and performance outcomes. However, there are some interesting differences between the third and fourth-generation models, which we will further elaborate in Section. 5.
@@
  • We successfully devise the black-box imperceptible character attack against a highly sophisticated publicly-accessible LLM system. In particularly, we present the four categories of methodologies to launch such adversarial attack for code analysis and comprehension task.
  • Extensive experiments are conducted to demonstrate the effectiveness of the attacks, covering a wide range of dataset and several state-of-the-art LLMs. We highlight the correlation between the attack budget and the performance outcomes.
@@ 2. Motivation
Recent advancements in machine learning technology have led to their widespread adoption across many different sectors. In particular, the ability to interact with a large language model using natural language has made them very easy to use for many people. However, as such systems have become more accessible – with ChatGPT exhibiting the fastest growth in userbase – the need to guarantee their security and reliability increases. Adversarial attacks against these models made through the prompting alone have been demonstrated to lead to a degradation in performance outcomes. There exists extant research confirming the deleterious effect of various attack techniques that inject perturbations into the prompt to manipulate the outcome of the LLMs. However, most research focuses on more typical natural language processing tasks such as sentiment analysis (et al., 2009) or text classification (et al., 2022).
The particular focus of this work is on the intersection between natural language processing and software testing. A very popular application of LLMs is in tools or processes that assist with computer programming, with LLMs demonstrating extraordinary potential in tasks such as code generation (Chen et al., 2021) and code comprehension (Fang et al., 2023). However, due to the interaction with these tools being through the medium of natural language, they are vulnerable to sophisticated attack vectors that target the discrepancy between human comprehension and software comprehension. For example, the Unicode specification (Consortium, 2021) defines many special Unicode characters that affect the rendering of symbols. This can lead to the content that the screen shows the user being extremely different to the content that the computer ‘sees’.
@@
Since most users will interact with LLM-integrated tools at the UI level rather than on a deeper, more technical level, it’s important to ensure awareness even robustness in the LLM against attacks that manipulate the UI level. Conducting a comprehensive investigation into these human-imperceptible attacks on large-language models has the aim of providing a better understanding of the internal mechanisms of the most popular models in current use, but also in identifying mitigation methods for the development of future generations of models.
What is displayed to a human can be significantly different to what is read by a large language model. There is an inherent difference in a user’s intent and the instructions given to an large language model. In the case where there are imperceptible perturbations made to a prompt, it is paramount to understand whether and to what extent an LLM can still perform its assigned tasks successfully. It helps in closing the gap between user interaction and system instructions. In this work, we investigate this challenging task by devising four categories of human-imperceptible attacks against different large language models, in particularly for three generations of large language models. We also explore the impact of advancements made in more recent generations on their robustness.
@@ 3.3. Attacks against LLMs
4. Study design
Figure 1. Experiment process diagram
The central aim of this work is to examine the effectiveness and identify the limitations of imperceptible-character-based attacks against LLMs oriented towards code comprehension tasks.
@@ 4.1. Research questions
To guide this investigation, and keeping in line with the overall research aims and objectives, this study seeks to answer the following questions:
@@
    To what extent do imperceptible perturbations within code correlate with LLM performance outcomes in software testing scenarios?
@@
    Does a correlation exist between imperceptible perturbations and LLM ‘confidence’, and what implications does this have for software testing?
@@
    How significantly do different imperceptible perturbation methods impact the responses of an LLM model?
@@
    How do advancements in recent LLM iterations impact their robustness against imperceptible adversarial perturbations in software testing scenarios?
@@ 4.2. Methodology
This study tested three different GPT models. We framed the research using a unified prompting strategy, which consists of three key elements:
@@
Using the OpenAI Chat Completions API (cha, 2024), the LLM was prompted to answer that same question for different pairs of code snippet and description. Since this study only aimed to evaluate the effect of code perturbations and not natural language perturbations, only the code snippet was perturbed, and the description remained constant. The response was limited to a binary output, i.e. Yes/No, so that the ‘logprobs’ feature of the OpenAI API (log, 2024) could be incorporated during the data analysis stage to deduce response confidence accurately. The entirety of the response data generation was executed programmatically, with no human interaction necessary. The study was also designed in a way that does not require human review of LLM responses to determine correctness.
4.3. Models tested
4.4. Dataset
To find the dataset for this experiment, inspiration was taken from a similar study (Jalil et al., 2023) which also aimed to prompt ChatGPT with software-related questions. In that study, ChatGPT was tasked with answering questions from a ‘popular software testing textbook’. Since this study needed a greater number of code snippets, that solution was not considered, but a similar solution was arrived at.
A dataset of 2644 LeetCode coding questions was found through huggingface.co (hug, 2023), an AI community website featuring user-submitted models and datasets. This dataset was deemed appropriate due to containing code snippets as well as their related descriptions. This meant that it would not be necessary to find a ‘base truth’ for all code snippets due to the veracity of provided descriptions being effectively guaranteed by LeetCode. The questions in the dataset have three categories of difficulty: easy, medium, and hard. Each question has answers available in four different coding languages: java, c++, python, and JavaScript. For this experiment, only JavaScript code snippets were used, to control for the confounding variable of different languages.
@@ 4.5. Attack taxonomy
In this section, we discuss our solution for the imperceptible character attack design, which devises four categories of perturbation for evaluation, each featuring a special Unicode character implementation:
  • Reordering: Character U+202E is used so the Unicode specification can support characters from languages that read from right-to-left (202, 2023). Inserting it at the head and tail of the reversed string means that to the human eye, the perturbed section is rendered in the correct order, but the input will be the reversed string.
  • Invisible characters: Character U+200C is a character that does not render to a visible glyph. When placed between two characters that would otherwise be joined by a ligature, it keeps them disconnected. It is also effectively a space character.
  • Deletions: Character U+0008 is the backspace character. It can be injected into a string to remove characters from visual rendering.
  • Homoglyphs: Homoglyphs are unique character that render to the same glyph. The Unicode Consortium publishes a document with the Unicode Security Mechanisms technical report (int, 2023) which details a mapping of characters that are intended as homoglyphs within the Unicode specification. These characters are intended to be mapped to the same glyph in font implementations and are thus ‘true’ homoglyphs. The technical report also includes a document which details a set of characters that are commonly visually confused, although these glyphs are not intended to be mapped to the same glyph in font implementations (uni, 2023). Since this second set is dependent on the choice of font implementation, and the extent to which they are visually similar is subjective, they were not considered for this study.
@@ 4.6. Attack generation
    Each perturbation is rendered identically to the clean code
@@
    The perturbations are inserted starting from beginning to end
@@
    The levels of perturbation budget are set at 20%, 40%, 60%, 80%, and 100%
@@
    The perturbations are created with special Unicode characters, but not exclusively so
@@
    The floor is taken for any mathematical operation involving non-integer numbers
@@ 4.7. Prompting
All three GPT LLMs were accessed through the OpenAI API with the written script, which can be found in the replication package. Each prompt was made using the Chat Completions API. A zero-shot prompting technique was used. The template of each prompt remained consistent across all categories of perturbations and all models.
The prompt template consisted of the following three messages:
@@ 4.8. Logprobs
Since there was no variation in token length, that confounding variable was controlled against. In fact, this limitation was conceived after running a previous iteration of the study that did not limit the token length of the response. The data gathered from this initial run was extremely inconsistent, as the log probability of every token was affected by its context.It proved impossible to determine which tokens were significant and which were superfluous, so the limitation to keep the response one token long was implemented.
@@
    The confidence of the LLM in this response

### v2

@@ Abstract.
Recent studies have demonstrated outstanding capabilities of large language models (LLMs) in software engineering tasks, including code generation and comprehension. While LLMs have shown significant potential in assisting with coding, it is perceived that LLMs are vulnerable to adversarial attacks. In this paper, we investigate the vulnerability of LLMs to imperceptible attacks, where hidden character manipulation in source code misleads LLMs’ behaviour while remaining undetectable to human reviewers. We devise these attacks into four distinct categories and analyse their impacts on code analysis and comprehension tasks. These four types of imperceptible coding character attacks include coding reordering, invisible coding characters, code deletions, and code homoglyphs. To comprehensively benchmark the robustness of current LLMs solutions against the attacks, we present a systematic experimental evaluation on multiple state-of-the-art LLMs. Our experimental design introduces two key performance metrics, namely model confidence using log probabilities of response, and the response correctness. A set of controlled experiments are conducted using a large-scale perturbed and unperturbed code snippets as the primary prompt input. Our findings confirm the susceptibility of LLMs to imperceptible coding character attacks, while different LLMs present different negative correlations between perturbation magnitude and performance. These results highlight the urgent need for robust LLMs capable of manoeuvring behaviours under imperceptible adversarial conditions. We anticipate this work provides valuable insights for enhancing the security and trustworthiness of LLMs in software engineering applications.
Adversarial attack, Large language models, Code comprehension
conference: The 29th International Conference on Evaluation and Assessment in Software Engineering; 17–20 June, 2025; Istanbul, Türkiyeccs: Computing methodologies Artificial intelligenceccs: Software and its engineeringccs: Software and its engineering General programming languages
@@ 1. Introduction
The success of Large Language Models (LLMs) in natural language processing (NLP) has driven their widespread adoption across many different sectors, one of which is software development. The presented ability to understand and generate code snippets through LLMs makes them highly accessible, resulting in the integration into developers’ workflows, including code generation (Du et al., 2024), program repair (Zhang et al., 2023), and vulnerability detection (Zhou et al., 2024).
However, the human-like responses of LLMs do not constitute as true reasoning capabilities, and their integration into software engineering workflows introduces new security risks and vulnerabilities. Considering their increasing adoption and continued growth in their sector, it is important to ensure the robustness of LLMs against tempted adversarial attacks and study their response mechanisms. Understanding these weaknesses in LLMs is fundamental to the future LLM-based software development.
@@
Recently, there are several studies demonstrating the feasibility of text-based adversarial attacks against LLMs, targeting performance outcomes of NLP tasks (Boucher et al., 2022; Formento et al., 2021; Li et al., 2020). While these studies have successful featured the LLMs vulnerabilities, it remains unknown concerning the feasibility of these attacks in real-world scenarios.  (Dyrmishi et al., 2023) argues that for an adversarial attack to be practical against a text-based LLM, the generated text must be both valid and natural. For attacks that fail to maintain high semantic consistency or the adversarial examples appear artificially generated, the corresponding real-world effectiveness will be limited. Through a survey of 378 human participants, their study concludes that existing text-based adversarial attacks lack imperceptibility, making them impractical in real-world scenarios where human reviewers are involved.
To address this research gap, this paper presents novel imperceptible adversarial attacks on source code, specifically targeting LLMs performing code analysis and comprehension tasks. Our goal is to understand whether or not the LLMs’ behaviour will be altered given manipulated code snippet which is visually indistinguishable from its original version when viewed by a human. We define ‘true imperceptibility’ as a malicious crafted code snippet that is completely indistinguishable from its original form. Prior works, such as (Liu et al., 2021), introduce imperceptible attacks in NLP but fail to achieve true imperceptibility, as the manipulated prompts, while semantically and visually similar, are still detectable when placed side-by-side. Our work bridges the gap by evaluating the practicality and effectiveness of truly imperceptible character attacks in real-world coding tasks with LLMs.
@@
Thus, in this work, our attack method introduces a group of perturbation methods in code snippets to maintain true imperceptibility. Specifically, we leverage the special Unicode specification characters that subtly affect visual rendering while preserving readability for human reviewers. Inspired by Boucher et al. (Boucher et al., 2022), we develop four distinct imperceptible attack strategies targeting code analysis and comprehension tasks, namely code reordering, invisible coding characters, code deletions, and code homoglyphs. We include different LLMs solutions for the attack performance evaluation. The efficacy of these attacks are assessed based on two key performance metrics in a large-scale code analysis task, which are model confidence and response correctness. Experiment results demonstrates a strong negative correlation between perturbation and performance outcomes, with some notable differences among these LLMs, which we will further elaborate in Section. 5.
@@
  • We investigate the vulnerability of LLMs to imperceptible adversarial attacks in the context of code analysis and comprehension tasks with LLMs.
  • We propose a black-box imperceptible coding character attack framework that introduces four distinct attack strategies against the state-of-the-art LLMs solutions. The strategies includes code reordering, invisible coding characters, code deletions, and code homoglyphs, which effectively manipulate LLMs’ behavior while remaining visually imperceptible for code analysis and comprehension task.
  • Extensive experiments are conducted to demonstrate the effectiveness of the attacks, covering a wide range of dataset and LLMs. We highlight the correlation between the attack budget and the performance outcomes.
@@
The remaining sections of the paper are as follows. Section 2 presents the motivation of this study, while Section 3 discusses the background and related works. We provide the details of the overall framework design and experiment details in Section 4, and the experimental results in Section 5. Section 6 discusses our findings with novel insights, and Section 7 summarizes the paper.
@@ 2. Motivation
LLMs have recently transformed different sectors with enhanced accessibility through natural language. However, with LLMs being increasingly integrated into critical applications, concerns about their security, robustness and vulnerability have grown. Adversarial attacks via prompt injection have been demonstrated to degrade LLMs’ performance. While existing works have investigated the vulnerability on LLMs, they primarily focus on typical NLP tasks, such as sentiment analysis (et al., 2009) or text classification (et al., 2022).
This work is motivated by the intersection of software testing and NLP, particularly for LLM assisted software development tasks. LLMs have demonstrated exceptional capabilities in tasks such as code generation (Chen et al., 2021) and code comprehension (Fang et al., 2023). However, their behaviour heavily rely on the prompts, which introduces novel vulnerabilities to attack vectors that exploit the discrepancy between human interpretation and LLMs’ comprehension. For example, there are different special characters that can manipulate text and symbol rendering from Unicode specification (Consortium, 2021), as depicted in Table 1. This may create a significant divergence between what is displayed to the user and what the computers ‘see’.
@@
As most users interact with LLM-integrated tools through the user interface (UI) rather than at a deeper technical level, ensuring robustness against attacks that manipulate UI-level representation is critical. Conducting a comprehensive investigation into these human-imperceptible attacks on large-language models has the aim of not only providing a better understanding of the internal mechanisms of LLMs in current use, but also in identifying mitigation methods for the development of future generations of LLMs.
Thus, we argue that what is displayed to a human can be significantly different to what the large language models ‘see’. These perturbations as introduced in the LLMs’ system present potential ways to alter the LLMs’ behaviour. To this end, it is crucial to assess whether and to what extent an LLM can still execute its assigned tasks correctly when imperceptible perturbations are leveraged. It helps in closing the gap between user interaction and system instructions. To this end, in this work, we systematically investigate this challenge by devising four categories of imperceptible coding character attacks against LLMs for code analysis task. We also explore the different performance presented in different LLMs.
@@ 3.3. Attacks against LLMs
4. Framework and Experiment Design
Figure 1. Overall imperceptible coding character attack framework
In this work, we aim to evaluate the effectiveness and identify the limitations of imperceptible coding character attacks on LLMs in the context of code comprehension tasks. Following we will outline the details of proposed research questions, overall black-box imperceptible coding character attack framework, and the experiment design.
@@ 4.1. Research questions
To guide the investigation, this study seeks to answer the following four questions:
@@
    To what extent do imperceptible perturbations within code correlate with LLMs’ performance outcomes in software testing scenarios?
@@
    Does a correlation exist between imperceptible perturbations and LLMs’ ‘confidence’, and what implications does this have for software testing?
@@
    How significantly do different imperceptible perturbation methods impact the responses of LLMs?
@@
    How do advancements in recent LLMs development improve their robustness against imperceptible adversarial perturbations in software testing scenarios?
@@ 4.2. Methodology
Figure 1 is the overall attack framework. Firstly, we frame the research using a unified prompting strategy in this work, which consists of three key elements:
@@
Using the OpenAI Chat Completions API (cha, 2024), the LLM is prompted to answer that same question for different pairs of code snippet and description. Since this study focuses on evaluating the effect of code perturbations rather than the natural language perturbations, only the code snippets are perturbed, and the descriptions remain constant. The response is limited to a binary output, i.e. Yes/No, so that we can leverage ‘logprobs’ (log, 2024) for accurately assessing response confidence during data analysis. The entire response generation process is fully automated, requiring no human interaction. Additionally, we have particularly designed the experiments that eliminate the need for human review to determine the correctness of LLM-generated responses. In next section, we will provide the details of the attack strategies.
@@ 4.3. Attack taxonomy
Herein, we discuss our solution for the imperceptible coding character attack design, which devises four categories of perturbation design, each featuring a special Unicode character implementation:
  • Code reordering: Character U+202E is used so the Unicode specification can support characters from languages that read from right-to-left (202, 2023). Inserting it at the head and tail of the reversed string means that to the human eye, the perturbed section is rendered in the correct order, but the input will be the reversed string.
  • Invisible coding characters: Character U+200C is a character that does not render to a visible glyph. When placed between two characters that would otherwise be joined by a ligature, it keeps them disconnected. It is also effectively a space character.
  • Code deletions: Character U+0008 is the backspace character. It can be injected into a string to remove characters from visual rendering.
  • Code homoglyphs: Homoglyphs are unique character that render to the same glyph. The Unicode Consortium publishes a document with the Unicode Security Mechanisms technical report (int, 2023) which details a mapping of characters that are intended as homoglyphs within the Unicode specification. These characters are intended to be mapped to the same glyph in font implementations and are thus ‘true’ homoglyphs. The technical report also includes a document which details a set of characters that are commonly visually confused, although these glyphs are not intended to be mapped to the same glyph in font implementations (uni, 2023). Since this second set is dependent on the choice of font implementation, and the extent to which they are visually similar is subjective, they were not considered for this study.
@@ 4.4. Attack generation
    Each perturbation is rendered identically to the clean code.
@@
    The perturbations are inserted starting from beginning to end.
@@
    The levels of perturbation budget are set at 20%, 40%, 60%, 80%, and 100%.
@@
    The perturbations are created with special Unicode characters, but not exclusively so.
@@
    The floor is taken for any mathematical operation involving non-integer numbers.
@@ 4.5. Prompting
We craft the prompts with the zero-shot prompting technique. We ensure that the templates of each prompt remain consistent across all categories of perturbations and all models. Following the implementation of different attack strategies, we leverage the Chat Completions API provided by the OpenAI API to access the tested LLMs, which can be found in the replication package. The prompt template consists of the following three messages:
@@ 4.6. Logprobs
Since there is no variation in token length, that confounding variable is controlled against. In fact, this limitation is conceived after running a previous iteration of the study that did not limit the token length of the response. The data gathered from this initial run is extremely inconsistent, as the log probability of every token is affected by its context.It proves impossible to determine which tokens are significant and which are superfluous, so the limitation to keep the response one token long is implemented.
@@
    The confidence of the LLM in this response (Conf.%).
4.7. Selected LLMs System
4.8. Dataset
To identify a large-scale dataset suitable for this study, inspiration is taken from a similar study (Jalil et al., 2023) which also aimed to prompt ChatGPT with software-related questions. In that study, ChatGPT is tasked with answering questions from a ‘popular software testing textbook’. Since this study needs a greater number of code snippets, that particular solution is not followed for implementation, but a similar solution is arrived at.
A dataset of 2644 LeetCode coding questions is found through huggingface.co (hug, 2023), an AI community website featuring user-submitted models and datasets. This dataset is deemed appropriate due to containing code snippets as well as their related descriptions. This means that it would not be necessary to find a ‘base truth’ for all code snippets due to the veracity of provided descriptions being effectively guaranteed by LeetCode. The questions in the dataset have three categories of difficulty: easy, medium, and hard. Each question has answers available in four different coding languages: java, c++, python, and JavaScript. For this experiment, only JavaScript code snippets are used, to control for the confounding variable of different languages.
@@ 5. Results
5.1. Overview
Figure 2. Correctness vs. perturbation budget for the Code Deletions attack across three models.
We investigated the robustness of three GPT-based language models, two GPT-3.5 variants and one GPT-4o version, against four code perturbation methods: reordering, invisible characters, deletions, and homoglyphs. Our findings show that all models consistently achieve 100% correctness on unaltered (clean) code, demonstrating their baseline ability to comprehend straightforward snippets. However, the GPT-3.5 models generally display a uniform, step-wise decline in confidence and correctness as the perturbation budget increases, particularly for deletions and reordering, while invisible characters and homoglyphs prove slightly less disruptive. In contrast, GPT-4o exhibits a drastic performance drop as soon as any level of perturbation is introduced, yielding near-zero correctness in most cases and indicating that the newer model may have additional filtering or checks that invalidate the perturbed code. These trends highlight how different perturbation types undermine model confidence at varying rates and underscore the stark contrast in failure modes between the GPT-3.5 and GPT-4o generations.